Find three consecutive integers such that the sum of the second and the third is 30 greater than the first
Answer by ikleyn(52874)

.
Find three consecutive integers such that the sum of the second and the third is 30 greater than the first
~~~~~~~~~~~~~~~~~~~~~

Let x be the smallest integer of the given three.
Then the integers are x, (x+1) and (x+2).


You are given that


(x+1) + (x+2) = x + 30.


Simplify:


2x + 3 = x + 30  --->  x = 30-3 = 27.


Answer.  The integers are 27, 28 and 29.